Circular No. 5880

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SGR 1806–20 Y. Tanaka, Institute of Space and Astronautical Science, on behalf of

the ASCA Team, communicates: “Following the notice of new activity of SGR 1806–20 (IAUC 5875), we observed the source with ASCA between Oct. 9.67 and 11.58 UT. Quick-look data show an anomalous increase of counts (burst) in a 4-s time bin, which includes the time of occurrence of a γ-ray burst (at Oct. 9.934) detected by the BATSE instrument on the Compton Gamma Ray Observatory from SGR 1806–20 (private communication by C. Kouveliotou). We discovered in the same observation a new x-ray source at α = 18^{h}05^{m}42^{s}, δ = −20^{o}25′10′′ (equinox 1950.0), with an error-circle radius of 1′, coincident with a radio source reported on IAUC 5879. The source intensity is ∼ 0.2 mCrab (0.5-10 keV) with a heavy absorption. Preliminary analysis localized the position of the ASCA burst into a cell of 3′ × 6′, which includes the new x-ray source. We tentatively identify this radio/x-ray source to be the counterpart of SGR 1806–20.”

V1974 CYGNI J. A. DeYoung and R. E. Schmidt, U.S. Naval Observatory, report:

“Time-resolved Kron-Cousins I-band photometry of V1974 Cyg (N Cyg 1992) on Oct. 6.99-7.08, 7.98-8.15, 10.98-11.15, and 12.98-13.18 UT shows a modulation with a period of 1.95 hr and a peak-to-peak amplitude of 0.16 ± 0.02 magnitude. A linear least-squares fit using the seven times of observed minima gives a period of 0.08123 ± 0.00003 day.”

NOTICE TO SUBSCRIBERS Subscribers to the Computer Service who have e-mail access are ad-

vised that, in addition to the login information they already have, use of the Username IAUMAIL (with the password information they already have) will allow them to MAIL to themselves specific Circulars that may have failed to arrive in the original e-mail transmission.

NOVA AQUILAE 1993 Further CCD photometry by H. Mikuž, Ljubljana, Slovenia (cf. IAUC

5818, 5838): Aug. 21.91 UT, V = 13.93 ± 0.05; Oct. 4.81, 13.29 ± 0.05.

1993 October 15 Daniel W. E. Green